# Minimax and adaptive estimation of general linear functionals under sparsity

### v1

@@ 6.2 Linear hypothesis testing
The proof of the lower bound in Theorem 8 follows arguments similar to those in Theorem 1; for completeness, the details are provided in Section C.3. For the upper bound, recall that the estimator $\hat{L}_{s}$ defined in (7) achieves the optimal squared error rate $\sigma^{2}\Phi_{\rm o}(s;\eta)$ over $\Theta_{s}$. Using this estimator, we construct the test
@@
for some constant $B>0$, and we show in Section A.6 that $\Delta_{s}$ achieves the minimax rate for testing with an appropriate choice of $B$. The proof also reveals that the results extend straightforwardly to the broader class of symmetric sub-Weibull noise distributions.
@@ 7 Future Directions
The sequence model in Equation 1 is closely related to the linear regression model, and our results yield new insights into the estimation of linear functionals in high-dimensional sparse regression. In Example 2, for orthogonal designs (i.e., $\mathbf{X}^{\top}\mathbf{X}$ is diagonal), the sparsity of $\theta$ corresponds directly to that of $\beta$. For high-dimensional linear regression, inference methods based on the idea of debiasing are well studied, but existing minimax theory relies on restrictive conditions on the loading vector $x_{0}$, such as bounded coordinate ratios or polynomial decay assumptions [2, 6]. These conditions may be relaxed since our work provides a sharp lower bound for minimax estimation with arbitrary loadings in the orthogonal design case, which may be extended to other cases. Another related and promising direction is to consider correlated noise, which has attracted considerable recent attention [19, 20].
@@ 8 Proof of the lower bounds
In this section, we establish the lower bounds stated in Theorems 1, 4 and 6. Without loss of generality, we set $\sigma=1$ in the proofs of the lower bounds. For two probability measures $P,Q$ defined on the same measurable space $(\mathcal{X},\mathcal{U})$, we define:
@@
$\mathrm{H}(P,Q)\;=\;\sqrt{\int\!\left(\sqrt{\frac{dP}{dQ}}-1\right)^{\!2}\,dQ},$
@@
$\chi^{2}(P\,\|\,Q)\;=\;\int\!\left(\frac{dP}{dQ}-1\right)^{\!2}dQ,$
@@ 8.1 General Tools
$\mathbb{P}_{\mu,\mathcal{P}_{\xi}}=\int_{\mathbb{R}^{d}}\mathbb{P}_{\theta,\mathcal{P}_{\xi}}\mu(d\theta).$
Specifically, we denote by $\mathbb{P}_{0,\mathcal{P}_{\xi}}$ the mixture probability measure when $\mu$ is the Dirac measure at 0, i.e., $\mathbb{P}_{0,\mathcal{P}_{\xi}}=\mathbb{P}_{\theta=0,\mathcal{P}_{\xi}}$. The following lemma, as a special case of [28, Theorem 2.15], is the key tool we will use to prove the lower bounds.
@@
Suppose that there exists two distributions $P_{1}=\mathbb{P}_{\mu_{1},\mathcal{P}_{\xi}^{1}}$ and $P_{2}=\mathbb{P}_{\mu_{2},\mathcal{P}_{\xi}^{2}}$ such that
@@
for some $c\in\mathbb{R},t>0,\beta_{1},\beta_{2},\beta_{3}\in(0,1)$, and some subset $\Theta\subseteq\mathbb{R}^{d}$. Then, for any estimator $\hat{T}$ of $L(\theta)$, we have
@@
$\int_{\mathbb{R}^{d}}\left[\sqrt{p(x;\theta)}-\sqrt{p(x;\theta+h)}\right]^{2}dx\leq\frac{|h|^{2}}{4}\int_{0}^{1}{\rm tr}\left[I(\theta+sh)\right]ds,\quad\forall\theta,h\in\mathbb{R}^{p}.$
@@ 8.2 Proof of the lower bounds in Theorem 1
We apply Lemma 1 with $\Theta=\Theta_{s}=\{\theta\in\mathbb{R}^{d}:\theta\in\Theta_{s}\}$, taking $P_{1}=\mathbb{P}_{0,F_{\alpha}^{\otimes}}$ and $P_{2}=\mathbb{P}_{\mu,F_{\alpha}^{\otimes}}$, where $\mu$ is a distribution on $\mathbb{R}^{d}$ that will be specified later. In what follows, we establish the lower bound for the case where $s>C_{1}$ and $\lambda_{o}s+\nu>C_{2}|\eta_{1}|$ with constants $C_{1},C_{2}>0$ to be specified. The proof for the complementary case is provided in the supplementary material.
@@
$\mu\left(\|\theta\|_{0}>s\right)\leq\mu\left(\|\theta\|_{0}-\mathbb{E}_{\mu_{1}}\|\theta\|_{0}>(1-\frac{c_{1}}{2})s\right)\leq\frac{{\rm Var}_{\mu_{1}}[\left\|\theta\right\|_{0}]}{(1-c_{1}/2)^{2}s^{2}}\leq\frac{c_{1}/2}{(1-c_{1}/2)^{2}s},$
@@
$\mu\left(L(\theta)<\frac{c_{1}}{8}C_{\alpha,2}[\lambda_{o}s+\nu]\right)$ $\leq\mu\left(L(\theta)<\frac{1}{2}\mathbb{E}_{\mu_{1}}L(\theta)\right)$
$\leq\frac{4{\rm Var}_{\mu_{1}}L(\theta)}{\left[\mathbb{E}_{\mu_{1}}[L(\theta)]\right]^{2}}$
@@
Let $P_{2}=\mathbb{P}_{\mu_{1},F_{\alpha}^{\otimes}}$, and denote by $f_{\alpha}^{j}$ the distribution of the $j$-th component under $P_{2}$. By Lemma 2, we have
@@ 8.3 Proof of Theorem 4
8.4 Proof of Theorem 6
$s^{2}\lambda^{2}_{\cal H}\geq C\sum_{j\leq s^{2}}\eta_{j}^{2}$
Note that for $j\geq s^{2}$, we have
and
Then, we have
$=$ $\exp(-s)\cdot\prod_{j=1}^{d}\left[1+\pi_{j}\cdot(e-1)\right]$
and
$\sigma_{j}=\sqrt{1-\gamma_{j}^{2}\cdot\pi_{j}(1-\pi_{j})}.$
Define $F_{0}=f_{0}^{\otimes d}$ and $F_{1}=\bigotimes_{i=1}^{d}f_{j}$. Therefore,
We bound the right hand side of the above inequality in two cases.
Let’s choose $\tau_{\alpha}=2/\log^{1/\alpha}2$.
$c\cdot\exp\left\{-\left(\frac{t}{2c_{2}}\right)^{\alpha}\right\}.$
Lemma 5.
This completes the proof of Equation 33 with Lemma 1.
Acknowledgement
D. Huang was partially supported by NUS Start-up Grant A-0004824-00-0.
@@ A.3 Proof of Theorem 3
Term II. Since by definition $\hat{s}\leq s_{0}$, we have $s<\hat{s}\leq s_{0}$ when the indicator inside the expectation equal 1. In this case, Lemmas 7 and 8 can be applied to obtain the following for any $\theta\in\Theta_{s}$:
@@
where the first inequality is due to Hölder’s inequality, and we have applied Lemmas 7 and 8 in the last line. Therefore, we have $\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\left[(\hat{L}_{\hat{s}}-L(\theta))^{2}\mathbbm{1}_{\hat{s}>s}\right]\leq c\cdot\Phi_{\rm adp}(1;\eta)$ for some constant $c>0$. The proof is completed since we have $\Phi_{\rm adp}(1;\eta)\leq c_{2,0}\Phi_{\rm adp}(s;\eta)$ and $\sum_{i=1}^{\infty}i^{-3/2}\log(ei)<\infty$. ∎
@@ A.7 Proofs of the Lemmas
Term I. The random variables $\eta_{j}\xi_{j}\mathbbm{1}\left\{|\eta_{j}\xi_{j}|\leq{~}\tau\lambda_{*}(s)\right\}$ are independent, mean zero and and bounded by $K={~}\tau\lambda_{*}(s)$. Let $U_{1}=\{j_{2}(s)<j\leq j_{2}(s^{\prime}):j\notin S\}$. From the Bernstein’s inequality for bounded distributions [30, Theorem 2.8.4], we have
@@
Moreover, when $\alpha\geq 2$, the truncated variables $\xi_{j}\mathbbm{1}\{|\eta_{j}\xi_{j}|\leq\tau\lambda_{*}(s)\}$ are sub-Gaussian, and thus Hoeffding’s inequality [30, Theorem 2.6.3] yields the same probability bound.
@@
for any $\Delta>0$. We can choose $\Delta=\zeta\Phi_{\rm adp}(s^{\prime};\eta)/(144\log(es^{\prime}))$ such that the first term on the right hand side of (49) is bounded by $2(s^{\prime})^{-8}$. Now for the second term, we will use the Fuk-Nagaev inequality [26, page 78], which we state here for the reader’s convenience.
@@ Appendix C Additional proof of the lower bounds
C.1 Proof of Lemma 2
@@
$c_{\sigma}=\int_{\mathbb{R}}\exp\!\left(-\frac{|x|^{\alpha}}{\sigma^{\alpha}}\right)\,dx=\sigma\!\int_{\mathbb{R}}e^{-|y|^{\alpha}}\,dy=\sigma\cdot\frac{2}{\alpha}\,\Gamma\!\left(\frac{1}{\alpha}\right).$
@@
$\int_{\mathbb{R}}x^{2}e^{-|x|^{\alpha}/\sigma^{\alpha}}\,dx=\sigma^{3}\int_{\mathbb{R}}y^{2}e^{-|y|^{\alpha}}\,dy=\sigma^{3}\cdot\frac{2}{\alpha}\,\Gamma\!\left(\frac{3}{\alpha}\right),$
@@
$\int_{\mathbb{R}}e^{-|x|^{\alpha}/\sigma^{\alpha}}\,dx=\sigma\int_{\mathbb{R}}e^{-|y|^{\alpha}}\,dy=\sigma\cdot\frac{2}{\alpha}\,\Gamma\!\left(\frac{1}{\alpha}\right).$
@@
$\frac{1}{c_{\sigma}}\int_{\mathbb{R}}x^{2}e^{-|x|^{\alpha}/\sigma^{\alpha}}\,dx=1\quad\Longleftrightarrow\quad\int_{\mathbb{R}}x^{2}e^{-|x|^{\alpha}/\sigma^{\alpha}}\,dx=\int_{\mathbb{R}}e^{-|x|^{\alpha}/\sigma^{\alpha}}\,dx,$
@@
$\sigma_{\alpha}^{2}=\frac{\displaystyle\int_{\mathbb{R}}e^{-|y|^{\alpha}}\,dy}{\displaystyle\int_{\mathbb{R}}y^{2}e^{-|y|^{\alpha}}\,dy}=\frac{\frac{2}{\alpha}\Gamma(1/\alpha)}{\frac{2}{\alpha}\Gamma(3/\alpha)}=\frac{\Gamma(1/\alpha)}{\Gamma(3/\alpha)}.$
@@
$\mathbb{E}_{F_{\alpha}(\sigma_{\alpha})}\exp\left(\frac{|X|^{\alpha}}{\tau^{\alpha}}\right)=\int_{\mathbb{R}}\frac{1}{c_{\sigma_{\alpha}}}\exp\left[\left(\frac{1}{\tau^{\alpha}}-\frac{1}{\sigma_{\alpha}^{\alpha}}\right)|x|^{\alpha}\right]\leq 2.$ (50)
@@
$\chi^{2}(f_{\alpha}^{(1)}\|f_{\alpha}^{(0)})=\int_{\mathbb{R}}\frac{f_{\alpha}^{(1)}(x)}{f_{\alpha}^{(0)}(x)}\,dx$ $=\frac{1}{c_{\sigma_{\alpha}}}\int_{\mathbb{R}}\exp\left(\frac{|x|^{\alpha}}{\sigma_{\alpha}^{\alpha}}-\frac{2|x-\gamma|^{\alpha}}{\sigma_{\alpha}^{\alpha}}\right)\,dx.$
@@
    $\chi^{2}(f_{\alpha}^{(1)}\|f_{\alpha}^{(0)})$ $=\frac{1}{c_{\sigma_{\alpha}}}\int_{\mathbb{R}}\exp\left(\frac{2|x|^{\alpha}}{\sigma_{\alpha}^{\alpha}}-\frac{2|x-\gamma|^{\alpha}}{\sigma_{\alpha}^{\alpha}}\right)\exp\left(-\frac{|x|^{\alpha}}{\sigma_{\alpha}^{\alpha}}\right)\,dx.$
    $\leq\frac{1}{c_{\sigma_{\alpha}}}\int_{\mathbb{R}}\exp\left(\frac{2|\gamma|^{\alpha}}{\sigma_{\alpha}^{\alpha}}\right)\exp\left(-\frac{|x|^{\alpha}}{\sigma_{\alpha}^{\alpha}}\right)\,dx\leq\exp\left(\frac{2|\gamma|^{\alpha}}{\sigma_{\alpha}^{\alpha}}\right)$
@@
    $\chi^{2}(f_{\alpha}^{(1)}\|f_{\alpha}^{(0)})$ $=\frac{1}{c_{\sigma_{\alpha}}}\int_{\mathbb{R}}\exp\left(\frac{|x|^{\alpha}}{\sigma_{\alpha}^{\alpha}}-\frac{2|x-\gamma|^{\alpha}}{\sigma_{\alpha}^{\alpha}}\right)\,dx.$
    $\leq\frac{1}{c_{\sigma_{\alpha}}}\int_{\mathbb{R}}\exp\left(\frac{s^{1-\alpha}|\gamma|^{\alpha}}{\sigma_{\alpha}^{\alpha}}\right)\exp\left(-\frac{(2-(1-s)^{1-\alpha})|x|^{\alpha}}{\sigma_{\alpha}^{\alpha}}\right)\,dx$
    $\leq\exp\left(\frac{2s^{1-\alpha}|\gamma|^{\alpha}}{\sigma_{\alpha}^{\alpha}}\right)\cdot\frac{1}{c_{\sigma_{\alpha}}}\int_{\mathbb{R}}\exp\left(-\frac{(1-2(1-s)^{1-\alpha})|x|^{\alpha}}{\sigma_{\alpha}^{\alpha}}\right)\,dx$
@@
    $\int_{\mathbb{R}}\exp\left(-\frac{(1-2(1-s_{\alpha})^{1-\alpha})|x|^{\alpha}}{\sigma_{\alpha}^{\alpha}}\right)\,dx<\infty.$
@@
    $\chi^{2}(f_{\alpha}^{(1)}\|f_{\alpha}^{(0)})\leq\frac{1}{c_{\sigma_{\alpha}}}\int_{\mathbb{R}}\exp\left(-\frac{(1-2(1-s_{\alpha})^{1-\alpha})|x|^{\alpha}}{\sigma_{\alpha}^{\alpha}}\right)\,dx\cdot\exp\left(\frac{2s_{\alpha}^{1-\alpha}|\gamma|^{\alpha}}{\sigma_{\alpha}^{\alpha}}\right).$
@@
$I(\tau)=$ $\int_{\mathbb{R}}\frac{1}{c_{\sigma_{\alpha}}}\exp\left[\left(\frac{1}{\tau^{\alpha}}-\frac{1}{\sigma_{\alpha}^{\alpha}}\right)|x|^{\alpha}\right]dx$
$=$ $\frac{2}{c_{\sigma_{\alpha}}}\int_{0}^{\infty}e^{-a(\tau)x^{\alpha}}dx$
$=$ $\frac{2}{c_{\sigma_{\alpha}}}\frac{1}{\alpha}\int_{0}^{\infty}y^{1/\alpha-1}e^{-a(\tau)y}dy$
@@
C.2 Additional proof of the lower bounds in  Theorem 1

### v2

@@ 6.2 Linear hypothesis testing
The proof of the lower bound in Theorem 8 follows arguments similar to those in Theorem 1; for completeness, the details are provided in Section C.3 in the supplementary material. For the upper bound, recall that the estimator $\hat{L}_{s}$ defined in (7) achieves the optimal squared error rate $\sigma^{2}\Phi_{\rm o}(s;\eta)$ over $\Theta_{s}$. Using this estimator, we construct the test
@@
for some constant $B>0$, and we show in Section A.6 in the supplementary material that $\Delta_{s}$ achieves the minimax rate for testing with an appropriate choice of $B$. The proof also reveals that the results extend straightforwardly to the broader class of symmetric sub-Weibull noise distributions.
@@ 7 Future Directions
The sequence model in Equation 1 is closely related to the linear regression model, and our results yield new insights into the estimation of linear functionals in high-dimensional sparse regression. For high-dimensional linear regression, inference methods based on the idea of debiasing are well studied, but existing minimax theory relies on restrictive conditions on the loading vector $x_{0}$, such as bounded coordinate ratios or polynomial decay assumptions [2, 6]. Our work suggests these conditions may be relaxed since the sharp lower bounds developed here provide an improved benchmark for arbitrary loadings in orthogonal designs (see Example 2). Another related and promising direction is to consider correlated noise, which has attracted considerable recent attention [18, 19].
@@ 8 Proof of the lower bounds
In this section, we establish the lower bounds stated in Theorems 1, 4 and 5. Without loss of generality, we set $\sigma=1$ in the proofs of the lower bounds. For two probability measures $P,Q$ defined on the same measurable space $(\mathcal{X},\mathcal{U})$, we define:
@@
$\mathrm{H}(P,Q)\;=\;\sqrt{\int\!\left(\sqrt{\frac{{\rm d\,}P}{{\rm d\,}Q}}-1\right)^{\!2}\,{\rm d\,}Q},$
@@
$\chi^{2}(P\,\|\,Q)\;=\;\int\!\left(\frac{{\rm d\,}P}{{\rm d\,}Q}-1\right)^{\!2}{\rm d\,}Q,$
@@ 8.1 General Tools
$\mathbb{P}_{\mu,\mathcal{P}_{\xi}}=\int_{\mathbb{R}^{d}}\mathbb{P}_{\theta,\mathcal{P}_{\xi}}\,\mu({\rm d\,}\theta).$
Specifically, we denote by $\mathbb{P}_{0,\mathcal{P}_{\xi}}$ the mixture probability measure when $\mu$ is the Dirac measure at 0, i.e., $\mathbb{P}_{0,\mathcal{P}_{\xi}}=\mathbb{P}_{\theta=0,\mathcal{P}_{\xi}}$. The following lemma, as a special case of [27, Theorem 2.15], is the key tool we will use to prove the lower bounds.
@@
Suppose that there exist two distributions $P_{1}=\mathbb{P}_{\mu_{1},\mathcal{P}_{\xi}^{1}}$ and $P_{2}=\mathbb{P}_{\mu_{2},\mathcal{P}_{\xi}^{2}}$ such that
@@
for some $c\in\mathbb{R}$, $t>0$, and $\beta_{1},\beta_{2},\beta_{3}\in(0,1)$, and some subset $\Theta\subseteq\mathbb{R}^{d}$. Then, for any estimator $\hat{T}$ of $L(\theta)$, we have
@@
$\int_{\mathbb{R}^{d}}\left[\sqrt{p(x;\theta)}-\sqrt{p(x;\theta+h)}\right]^{2}\nu({\rm d\,}x)\leq\frac{|h|^{2}}{4}\int_{0}^{1}{\rm tr}\left[I(\theta+sh)\right]{\rm d\,}s,\quad\forall\theta,h\in\mathbb{R}^{p}.$
@@ 8.2 Proof of the lower bounds in Theorem 1
We apply Lemma 1 with $\Theta=\Theta_{s}$, taking $P_{1}=\mathbb{P}_{0,F_{\alpha}^{\otimes}}$ and $P_{2}=\mathbb{P}_{\mu,F_{\alpha}^{\otimes}}$, where $\mu$ is a distribution on $\mathbb{R}^{d}$ that will be specified later. In what follows, we establish the lower bound for the case where $s>C_{1}$ and $\lambda_{o}s+\nu>C_{2}|\eta_{1}|$ with constants $C_{1},C_{2}>0$ to be specified. The proof for the complementary case is provided in the supplementary material.
@@
$\mu\left(\|\theta\|_{0}>s\right)\leq\mu\left(\|\theta\|_{0}-\mathbb{E}_{\mu}\|\theta\|_{0}>(1-\frac{c_{1}}{2})s\right)\leq\frac{{\rm Var}_{\mu}[\left\|\theta\right\|_{0}]}{(1-c_{1}/2)^{2}s^{2}}\leq\frac{c_{1}/2}{(1-c_{1}/2)^{2}s},$
@@
$\mu\left(L(\theta)<\frac{c_{1}}{8}C_{\alpha,2}[\lambda_{o}s+\nu]\right)$ $\leq\mu\left(L(\theta)<\frac{1}{2}\mathbb{E}_{\mu}L(\theta)\right)$
$\leq\frac{4{\rm Var}_{\mu}L(\theta)}{\left[\mathbb{E}_{\mu}[L(\theta)]\right]^{2}}$
@@
Let $P_{2}=\mathbb{P}_{\mu,F_{\alpha}^{\otimes}}$, and denote by $f_{\alpha}^{j}$ the distribution of the $j$-th component under $P_{2}$. By Lemma 2, we have
@@ 8.3 Proof of Theorem 4
Appendix A provides detailed proofs of all upper bound results in the paper. Appendix B provides the proof for the uniqueness of the solution to Equation 5 and includes the proofs of the propositions on the rate characteristics. Appendix C includes the proofs of Theorem 5 and the lemmas used in proving the lower bounds Section 8. Appendix D contains the details for the examples in Section 5.
@@ A.3 Proof of Theorem 3
Term II. Since by definition $\hat{s}\leq s_{0}$, we have $s<\hat{s}\leq s_{0}$ when the indicator inside the expectation equal 1. In this case, Lemmas 6 and 7 can be applied to obtain the following for any $\theta\in\Theta_{s}$:
@@
where the first inequality is due to Hölder’s inequality, and we have applied Lemmas 6 and 7 in the last line. Therefore, we have $\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\left[(\hat{L}_{\hat{s}}-L(\theta))^{2}\mathbbm{1}_{\hat{s}>s}\right]\leq c\cdot\Phi_{\rm adp}(1;\eta)$ for some constant $c>0$. The proof is completed since we have $\Phi_{\rm adp}(1;\eta)\leq c_{2,0}\Phi_{\rm adp}(s;\eta)$ and $\sum_{i=1}^{\infty}i^{-3/2}\log(ei)<\infty$. ∎
@@ A.7 Proofs of the Lemmas
Term I. The random variables $\eta_{j}\xi_{j}\mathbbm{1}\left\{|\eta_{j}\xi_{j}|\leq{~}\tau\lambda_{*}(s)\right\}$ are independent, mean zero and and bounded by $K={~}\tau\lambda_{*}(s)$. Let $U_{1}=\{j_{2}(s)<j\leq j_{2}(s^{\prime}):j\notin S\}$. From the Bernstein’s inequality for bounded distributions [28, Theorem 2.8.4], we have
@@
Moreover, when $\alpha\geq 2$, the truncated variables $\xi_{j}\mathbbm{1}\{|\eta_{j}\xi_{j}|\leq\tau\lambda_{*}(s)\}$ are sub-Gaussian, and thus Hoeffding’s inequality [28, Theorem 2.6.3] yields the same probability bound.
@@
for any $\Delta>0$. We can choose $\Delta=\zeta\Phi_{\rm adp}(s^{\prime};\eta)/(144\log(es^{\prime}))$ such that the first term on the right hand side of (48) is bounded by $2(s^{\prime})^{-8}$. Now for the second term, we will use the Fuk-Nagaev inequality [25, page 78], which we state here for the reader’s convenience.
@@ Appendix C Additional proof of the lower bounds
C.1 Proof of Theorem 5
$s^{2}\lambda^{2}_{\cal H}\geq C\sum_{j\leq s^{2}}\eta_{j}^{2}$
Note that for $j\geq s^{2}$, we have
and
Then, we have
$=$ $\exp(-s)\cdot\prod_{j=1}^{d}\left[1+\pi_{j}\cdot(e-1)\right]$
and
$\sigma_{j}=\sqrt{1-\gamma_{j}^{2}\cdot\pi_{j}(1-\pi_{j})}.$
Define $F_{0}=f_{0}^{\otimes d}$ and $F_{1}=\bigotimes_{i=1}^{d}f_{j}$. Therefore,
We bound the right hand side of the above inequality in two cases.
Let’s choose $\tau_{\alpha}=2/\log^{1/\alpha}2$.
$c\cdot\exp\left\{-\left(\frac{t}{2c_{2}}\right)^{\alpha}\right\}.$
Lemma 11.
This completes the proof of Equation 30 with Lemma 1.
C.2 Proof of Lemma 2
@@
$c_{\sigma}=\int_{\mathbb{R}}\exp\!\left(-\frac{|x|^{\alpha}}{\sigma^{\alpha}}\right)\,{\rm d\,}x=\sigma\!\int_{\mathbb{R}}e^{-|y|^{\alpha}}\,{\rm d\,}y=\sigma\cdot\frac{2}{\alpha}\,\Gamma\!\left(\frac{1}{\alpha}\right).$
@@
$\int_{\mathbb{R}}x^{2}e^{-|x|^{\alpha}/\sigma^{\alpha}}\,{\rm d\,}x=\sigma^{3}\int_{\mathbb{R}}y^{2}e^{-|y|^{\alpha}}\,{\rm d\,}y=\sigma^{3}\cdot\frac{2}{\alpha}\,\Gamma\!\left(\frac{3}{\alpha}\right),$
@@
$\int_{\mathbb{R}}e^{-|x|^{\alpha}/\sigma^{\alpha}}\,{\rm d\,}x=\sigma\int_{\mathbb{R}}e^{-|y|^{\alpha}}\,{\rm d\,}y=\sigma\cdot\frac{2}{\alpha}\,\Gamma\!\left(\frac{1}{\alpha}\right).$
@@
$\frac{1}{c_{\sigma}}\int_{\mathbb{R}}x^{2}e^{-|x|^{\alpha}/\sigma^{\alpha}}\,{\rm d\,}x=1\quad\Longleftrightarrow\quad\int_{\mathbb{R}}x^{2}e^{-|x|^{\alpha}/\sigma^{\alpha}}\,{\rm d\,}x=\int_{\mathbb{R}}e^{-|x|^{\alpha}/\sigma^{\alpha}}\,{\rm d\,}x,$
@@
$\sigma_{\alpha}^{2}=\frac{\displaystyle\int_{\mathbb{R}}e^{-|y|^{\alpha}}\,{\rm d\,}y}{\displaystyle\int_{\mathbb{R}}y^{2}e^{-|y|^{\alpha}}\,{\rm d\,}y}=\frac{\frac{2}{\alpha}\Gamma(1/\alpha)}{\frac{2}{\alpha}\Gamma(3/\alpha)}=\frac{\Gamma(1/\alpha)}{\Gamma(3/\alpha)}.$
@@
$\mathbb{E}_{F_{\alpha}(\sigma_{\alpha})}\exp\left(\frac{|X|^{\alpha}}{\tau^{\alpha}}\right)=\int_{\mathbb{R}}\frac{1}{c_{\sigma_{\alpha}}}\exp\left[\left(\frac{1}{\tau^{\alpha}}-\frac{1}{\sigma_{\alpha}^{\alpha}}\right)|x|^{\alpha}\right]{\rm d\,}x\leq 2.$ (50)
@@
$\chi^{2}(f_{\alpha}^{(1)}\|f_{\alpha}^{(0)})=\int_{\mathbb{R}}\frac{f_{\alpha}^{(1)}(x)}{f_{\alpha}^{(0)}(x)}\,{\rm d\,}x$ $=\frac{1}{c_{\sigma_{\alpha}}}\int_{\mathbb{R}}\exp\left(\frac{|x|^{\alpha}}{\sigma_{\alpha}^{\alpha}}-\frac{2|x-\gamma|^{\alpha}}{\sigma_{\alpha}^{\alpha}}\right)\,{\rm d\,}x.$
@@
    $\chi^{2}(f_{\alpha}^{(1)}\|f_{\alpha}^{(0)})$ $=\frac{1}{c_{\sigma_{\alpha}}}\int_{\mathbb{R}}\exp\left(\frac{2|x|^{\alpha}}{\sigma_{\alpha}^{\alpha}}-\frac{2|x-\gamma|^{\alpha}}{\sigma_{\alpha}^{\alpha}}\right)\exp\left(-\frac{|x|^{\alpha}}{\sigma_{\alpha}^{\alpha}}\right)\,{\rm d\,}x.$
    $\leq\frac{1}{c_{\sigma_{\alpha}}}\int_{\mathbb{R}}\exp\left(\frac{2|\gamma|^{\alpha}}{\sigma_{\alpha}^{\alpha}}\right)\exp\left(-\frac{|x|^{\alpha}}{\sigma_{\alpha}^{\alpha}}\right)\,{\rm d\,}x\leq\exp\left(\frac{2|\gamma|^{\alpha}}{\sigma_{\alpha}^{\alpha}}\right)$
@@
    $\chi^{2}(f_{\alpha}^{(1)}\|f_{\alpha}^{(0)})$ $=\frac{1}{c_{\sigma_{\alpha}}}\int_{\mathbb{R}}\exp\left(\frac{|x|^{\alpha}}{\sigma_{\alpha}^{\alpha}}-\frac{2|x-\gamma|^{\alpha}}{\sigma_{\alpha}^{\alpha}}\right)\,{\rm d\,}x.$
    $\leq\frac{1}{c_{\sigma_{\alpha}}}\int_{\mathbb{R}}\exp\left(\frac{s^{1-\alpha}|\gamma|^{\alpha}}{\sigma_{\alpha}^{\alpha}}\right)\exp\left(-\frac{(2-(1-s)^{1-\alpha})|x|^{\alpha}}{\sigma_{\alpha}^{\alpha}}\right)\,{\rm d\,}x$
    $\leq\exp\left(\frac{2s^{1-\alpha}|\gamma|^{\alpha}}{\sigma_{\alpha}^{\alpha}}\right)\cdot\frac{1}{c_{\sigma_{\alpha}}}\int_{\mathbb{R}}\exp\left(-\frac{(1-2(1-s)^{1-\alpha})|x|^{\alpha}}{\sigma_{\alpha}^{\alpha}}\right)\,{\rm d\,}x$
@@
    $\int_{\mathbb{R}}\exp\left(-\frac{(1-2(1-s_{\alpha})^{1-\alpha})|x|^{\alpha}}{\sigma_{\alpha}^{\alpha}}\right)\,{\rm d\,}x<\infty.$
@@
    $\chi^{2}(f_{\alpha}^{(1)}\|f_{\alpha}^{(0)})\leq\frac{1}{c_{\sigma_{\alpha}}}\int_{\mathbb{R}}\exp\left(-\frac{(1-2(1-s_{\alpha})^{1-\alpha})|x|^{\alpha}}{\sigma_{\alpha}^{\alpha}}\right)\,{\rm d\,}x\cdot\exp\left(\frac{2s_{\alpha}^{1-\alpha}|\gamma|^{\alpha}}{\sigma_{\alpha}^{\alpha}}\right).$
@@
$I(\tau)=$ $\int_{\mathbb{R}}\frac{1}{c_{\sigma_{\alpha}}}\exp\left[\left(\frac{1}{\tau^{\alpha}}-\frac{1}{\sigma_{\alpha}^{\alpha}}\right)|x|^{\alpha}\right]{\rm d\,}x$
$=$ $\frac{2}{c_{\sigma_{\alpha}}}\int_{0}^{\infty}e^{-a(\tau)x^{\alpha}}{\rm d\,}x$
$=$ $\frac{2}{c_{\sigma_{\alpha}}}\frac{1}{\alpha}\int_{0}^{\infty}y^{1/\alpha-1}e^{-a(\tau)y}{\rm d\,}y$
@@
C.3 Additional proof of the lower bounds in  Theorem 1